# Multi-Modal Diffusion for Hand-Object Grasp Generation

### v1

@@ Abstract
In this work, we focus on generating hand grasp over objects. Compared to previous works of generating hand poses with a given object, we aim to allow the generalization of both hand and object shapes by a single model. Our proposed method Multi-modal Grasp Diffusion (MGD) learns the prior and conditional posterior distribution of both modalities from heterogeneous data sources. Therefore it relieves the limitation of hand-object grasp datasets by leveraging the large-scale 3D object datasets. According to both qualitative and quantitative experiments, both conditional and unconditional generation of hand grasp achieve good visual plausibility and diversity. The proposed method also generalizes well to unseen object shapes. The code and weights will be available at https://github.com/noahcao/mgd
Figure 1: Applications of our proposed method Multi-modal Grasp Diffusion (MGD). Left: generating the hand grasp on an unseen object. Middle: jointly sampling a combination of the hand and the object for a plausible grasp. Right: the generated grasp can be used as guidance for generating photo-realistic grasping images using existing image generation tools [2].
 [^1] The work in this paper was done by Jinkun Cao during his internship at Adobe in 2023 summer
@@ 1 Introduction
In this paper, we focus on the synthesis of static hand grasps. Instead of fitting to a small set of objects presented in certain 3D hand-object interaction datasets, we wish to generalize to diverse object geometries. We build a multi-modal generative model, capable of generating the hand poses either conditional to or jointly with the object shapes. Our model can learn to grasp objects in a pure geometric perspective instead of requiring additional conditions such as language descriptions or category labels.
Existing works [26, 14, 7] for hand-object grasp generation typically relies on datasets with full-stack 3D annotations [26, 7, 33]. Moreover, existing datasets are designed for different purposes in animation and robotics. Their grasp annotations are constrained to certain hand parametric models [16, 12, 1, 30, 23, 26, 33], which prevents combining different data resources. Compared to the hand model [30] for robotic manipulation, MANO [23] is a non-rigid hand model with blend skinning, requiring more details, flexibility and degree of freedom. Thus, though large-scale synthetic datasets [30] are available for certain hand models, the datasets compatible with the MANO parametric model still suffer from limited scale and availability due to the difficulty of scanning and multi-view captures. On the other hand, human beings can transfer a grasping pattern to different objects and make plausible grasping choices for an unseen or even imagined object based on the object’s geometry. We call such ability universal grasp generation. Though full-stack 3D annotation is limited, we have large-scale 3D object datasets available. Therefore, in this work, we aim to study if we can combine the large-scale 3D object datasets to help hand-object grasp generation by jointly modeling hand and object representations in a latent space. Specifically, our model can accept grasping data with partial supervision in training, such as hand pose and object shape whichever are available, which significantly increases the datasets it can use for learning. Our model can generate either the full 3D grasping scenes with both the posed hand and object or a posed hand conditioned on a given object geometry. We believe such a flexible hand-object interaction generation ability without domain-specific auxiliary information makes one step forward for human-like universal grasp generation.
@@
Following such an intention, we propose our Multi-modal Grasp Diffusion (MGD) model. It follows the latent diffusion model [22] (LDM) to encode and ensemble different modalities into a latent space and then generate plausible samples by the probabilistic denoising process [11]. To learn from more diverse objects, we construct the model by leveraging the object shape encoding and decoding networks pre-trained on the large-scale generic 3D object shape dataset [4]. As we aim to derive a universal grasp generation solution, we remove the category-specific information when constructing the latent code for objects and encode the object shapes in a purely geometric fashion.
@@
As the hand articulation and positions are always coupled to the object shape in hand-object interaction, we design the latent diffusion to denoise the latent code from different modalities as a whole. However, we also need to decouple each modality’s representation to allow the model training by fusing data from different resources with heterogeneous annotations. Therefore, our method has independent encoder and decoder networks for different modalities and can optimize the generation of each modality solely. Such a design solves the limitation of object diversity in hand-object interaction datasets as we could leverage the data from generic object shape datasets to tune the object generation part. To achieve a certain disentanglement of modality representations, we propose to use asynchronous denoising schedulers for different modalities during noise diffusion and denoising in training. We believe we are the first to use such a strategy to solve the training stability problem raised by data with unpaired and heterogeneous annotations.
@@
To conclude, we develop a multi-modal hand grasp generation model to sample from either joint distribution for both hand and object or the object-conditioned distribution for hand generation. It has the advantage of incorporating training resources with different annotations. By the qualitative and quantitative evaluations, our proposed method shows a good performance for grasp generation in both unconditional and object-conditioned settings. Thanks to the posterior generation learned from rich data resources, it shows significantly better performance in generating grasp over out-of-domain object shapes, which is critical for universal grasp generation. Considering that no published works have supported the generation of both hand and object to form a grasp in an end-to-end fashion, we believe our work is pioneering for the universal grasp generation task.
@@
With the flexibility to generate both modalities in hand grasp, our model can facilitate many downstream applications. The grasps are good references to generate photo-realistic images with good geometry alignment and significantly fewer artifacts. We first generate grasps in 3D and use them as conditional signals for image generation can help produce high-quality hand images with challenging gestures and view angles. We demonstrate using grasp from our model as the condition improves grasp alignment on images and relieves artifacts. Some examples of object-conditioned grasp generation, joint generation, and image generation with grasp rendering are presented in Figure 1. The main contribution of this work is to propose a generative model capable of generating hand grasp in either unconditional or object-conditioned fashions and the corresponding strategy to enhance the performance with limited available 3D hand-object grasp annotations.
@@ 2 Related Works
Hand Grasp Generation. Modern generative models [11] are recently introduced into hand grasp reconstruction [34] and generation. Hand grasp reconstruction asks for hand grasp rendering aligned with images while hand grasp generation asks for hand grasp visually or physically plausible. GrabNet [26] is a commonly used method for hand grasp generation but it can hardly be generalized to diverse object shapes. HOIDiffusion [38] uses the 3D hand grasp from GrabNet and an image diffusion model to generate hand-object interaction images. AffordanceDiffusion [35] also studies the generation of hand-object interaction for image synthesis. Instead of generating hand-object interaction images or videos, we focus on improving hand-object interaction in 3D space. Many recent works are based on diffusion models [36, 14, 5] or physics-based simulators [37, 15, 31]. Compared to existing works which typically require a given object [26, 38, 35, 39] shape as input, we desire generating a hand grasp over a given object or generating both the object and hand to form a plausible grasp.
Multi-Modal Generation. With the rise of diffusion models, multi-modal generation has been extended in many areas, such as image+text generation [3] and audio+pixel generation [24]. Recently, a concurrent work UGG [14] studies to generate both hand and object to form a grasp but it uses the ShadowHand parametric model to leverage the large-scale synthetic ShadowHand grasping datasets [30] and the final results are optimization-based instead of directly from the generative model. On the other hand, another concurrent work G-HOP [36] builds a multi-modal hand-object grasp prior by encoding object shape and hand poses into a unified latent space but a language prompt is required to provide necessary conditional information. In this work, we focus on building a multi-modal diffusion model to generate the modalities of a 3D object and hand at the same time without requiring optimization or auxiliary conditions. The hand grasp is purely geometric-based. It allows more flexibility for downstream tasks, such as generating photo-realistic images by using the rendering of generated 3D hand-object interaction as a condition.
Figure 2: The illustration of our proposed Multi-modal Grasp Diffusion model (MGD). We present the main modules of the model while removing the secondary modules for simplicity. The optional object condition can turn the generation into object-conditioned.
@@ 3 Method
In  Section 3.1, we first provide the formal problem formulation. We then review the related preliminary knowledge for methodology and data representation in Section 3.2. Finally, we introduced our proposed method in Section 3.3.
@@ 3.1 Problem Formulation
An instance of hand-object interaction consists of a posed hand, denoted as $\mathbf{H}$, and a posed object, denoted as $\mathbf{O}$. To learn the relationship between hand gestures and object shapes, we model the joint distribution of hand and object:
$\vspace{-0.1em}\{\mathbf{O},\mathbf{H}\}\sim\mathcal{D}_{\Phi}.\vspace{-0.1em}$ (1)
Also, in the community of visual perception, animation, and robotics, people are also interested in generating hand grasp on certain objects, which can be formulated as
$\vspace{-0.1em}\{\mathbf{H}\}\sim\mathcal{D}_{\Phi}|_{\mathbf{O}_{g}},\vspace{-0.1em}$ (2)
where $\mathbf{O}_{g}$ is a provided object shape. In the previous works, the unconditional generation has been under-explored. And stand-only models for conditional hand grasp generation are usually required with no flexibility for object shape generation. Stand-only models suffer from bad robustness to different object shapes and grasp diversity. This is because existing datasets with 3D annotations of both modalities are notoriously limited in covered object shapes due to the expansiveness of scanning the object shape and capturing the corresponding hand poses. In contrast, we aim to solve both scenarios by a single model. During training, this model can adapt to any supervision which maximizes the utilization of available data.
@@ 3.2.1 Latent Diffusion
Our method follows the latent diffusion models (LDM) [22, 25] to generate samples in two modality spaces, i.e., articulable hands, and rigid objects. LDM diffuses and denoises in a latent space instead of on the raw data representations.
@@
where $\alpha_{t}\in(0,1)$ are constant hyper-parameters determined by the time step $t$ and a noise scheduler. Though the typical diffusion model [11] trains a denoiser to predict the noise added per time step during the diffusion stage, there is another line of diffusion models [21, 28] that learns to recover the clean data sample directly. The loss turns to
@@ 3.2.2 Modality Representations
Hand. We follow the MANO [23] parametric model to represent hands. Compared to another parametric model ShadowHand [1], MANO is designed for animating non-rigid real human hands instead of robotic hands thus a better fit for animation and photo-realistic generalization tasks. However, its higher complexity causes difficulty in grasping data synthesis. Therefore, there is no MANO-based hand grasping dataset as large-scale as the synthetic ShadowHand-based datasets [30] yet. This prevents research on the joint generation of MANO hands and objects in a straightforward supervised learning fashion. In this work, we directly use MANO parameters as joint rotation angles, i.e. $\theta\in\mathbb{R}^{48}$, instead of the joint xyz position to avoid the inverse kinematics. We also use the translation of hands $\mathbf{t}_{h}\in\mathbb{R}^{3}$. We don’t stylize the hand shapes, thus we use constant mean shape parameters $\beta_{0}=\mathbf{0}^{10}$ during linear blend skinning. The parameters of hands are concatenated and encoded by a hand encoder $\mathcal{E}_{H}$ to be the hand latent code $\mathbf{y}_{H}\in\mathbb{R}^{51}$.
Object. For object representation, we use the point cloud latent code as $\mathbf{y}_{O}=\mathbf{h}\in\mathbb{R}^{N\times 4}$, where $N$ is the number of points. As the grasp is modeled in the object-centric frame, no translation or orientation is required in the object pose and shape representation anymore. This convention is aligned with the definition of data in many hand-object grasp datasets, such as OakInk [33].
@@ 3.2.3 Object Generation
The limited amount and high category-specific bias of the objects in hand-object interaction (HOI) datasets prevent learning robust and generalizable grasps. By leveraging the large-scale object-only 3D datasets, we can derive a more robust object encoding for the grasp generation than learning from the HOI datasets solely. For the object encoding and generation part, we borrow the point-cloud-based object shape generation method LION [29]. Existing HOI generation methods can only learn object shapes from HOI datasets, including only dozens of objects, while LION learns from a much larger basis, i.e., more than 50,000 objects in ShapeNet [4]. LION proposes a Hierarchical Latent Point Diffusion Model to generate objects in two stages. In the first step, it derives a global latent code $\mathbf{z}^{G}_{0}\in\mathbb{R}^{D_{G}}$ from a posterior distribution $q_{\phi}(\mathbf{z}^{G}_{0}|\mathbf{P})$ where $\mathbf{P}\in\mathbb{R}^{u\times 3}$ is the object point cloud. Then it samples a point cloud latent $\mathbf{h}_{0}$ from a posterior $q_{\mathcal{E}_{O}}(\mathbf{h}_{0}|\mathbf{P},\mathbf{z}^{G}_{0})$. LION generates the object shape by a reverse process from sampled latent code
$\vspace{0cm}p_{\mathcal{E}_{O},\psi,\gamma}(\mathbf{P},\mathbf{h}_{0},\mathbf{z}^{G}_{0})=p_{\mathcal{X}_{O}}(\mathbf{P}|\mathbf{h}_{0},\mathbf{z}^{G}_{0})p_{\psi}(\mathbf{h}_{0}|\mathbf{z}^{G}_{0})p_{\gamma}(\mathbf{z}^{G}_{0}),\vspace{0cm}$ (5)
@@
$\vspace{0cm}p_{\mathcal{E}_{O},\psi^{\prime}}(\mathbf{P},\mathbf{h}_{0})=p_{\mathcal{X}_{O}}(\mathbf{P}|\mathbf{h}_{0})p_{\psi^{\prime}}(\mathbf{h}_{0}),\vspace{0cm}$ (6)
where $\psi^{\prime}$ is fine-tuned from $\psi$ by replacing the global prior $\gamma$ with a standard Gaussian. Thanks to the pre-training from large-scale object datasets [4], we could use the weights of $\mathcal{E}_{O}$ and $\mathcal{X}_{O}$ from LION directly. By freezing the encoder and decoder, we fine-tune the object prior distribution with the object shapes from HOI datasets. In this fashion, the model provides a robust object encoding for the latent diffusion model and keeps the sampled object shapes proper for the single-hand interaction.
3.3 Multi-modal Grasp Diffusion (MGD)
We now introduce our proposed Multi-modal Grasp Diffusion (MGD). We present the overall illustration and key designs of our proposed method in Figure 2.

### v2

# Joint Diffusion for Universal Hand-Object Grasp Generation

@@ Abstract
Predicting and generating human hand grasp over objects is critical for animation and robotic tasks. In this work, we focus on generating both the hand and objects in a grasp by a single diffusion model. Our proposed Joint Hand-Object Diffusion (JHOD) models the hand and object in a unified latent representation. It uses the hand-object grasping data to learn to accommodate hand and object to form plausible grasps. Also, to enforce the generalizability over diverse object shapes, it leverages large-scale object datasets to learn an inclusive object latent embedding. With or without a given object as an optional condition, the diffusion model can generate grasps unconditionally or conditional to the object. Compared to the usual practice of learning object-conditioned grasp generation from only hand-object grasp data, our method benefits from more diverse object data used for training to handle grasp generation more universally. According to both qualitative and quantitative experiments, both conditional and unconditional generation of hand grasp achieves good visual plausibility and diversity. With the extra inclusiveness of object representation learned from large-scale object datasets, the proposed method generalizes well to unseen object shapes.
*: Jinkun and Yi were affilated with Adobe Research when the main part of the work was conducted
@@ 1 Introduction
In this paper, we explore the generation of hand-object grasps. Unlike traditional approaches that fit to a limited set of objects found in 3D hand-object interaction datasets, our goal is to generalize across diverse object geometries. We develop a joint diffusion model capable of generating hand poses either conditionally on or jointly with object shapes. Our method approaches objects from a purely geometric perspective, eliminating the need for language descriptions or category labels. Despite having access to limited hand-object interaction datasets, our model learns a comprehensive object shape embedding by leveraging large-scale object shape datasets. It generates paired hand and object grasps by denoising a joint latent representation.
Figure 1: Applications of our proposed method Joint Hand-Object Diffusion (JHOD). Left: generating the hand grasp on an unseen object. Middle: jointly sampling a combination of the hand and the object for a plausible grasp. Right: the generated grasp can be used as guidance for generating photo-realistic grasping images using existing image generation tools (Adobe, 2024).
To generate hand grasp over objects, existing works (Taheri et al., 2020; Lu et al., 2023; Fan et al., 2023) typically rely on datasets with full-stack 3D annotations (Taheri et al., 2020; Fan et al., 2023; Yang et al., 2022) from object scans and hand pose capture. However, with the difficulty of capturing 3D object models and hand gestures, this area of research faces notorious limitations of annotated data,e.g., only dozens of object shapes are available in a dataset (Fan et al., 2023; Taheri et al., 2020). Moreover, existing datasets are designed for different hand parametric models (A. T. Miller and P. K. Allen (2004); J. Jian, X. Liu, M. Li, R. Hu, and J. Liu (2023); 6; R. Wang, J. Zhang, J. Chen, Y. Xu, P. Li, T. Liu, and H. Wang (2023); J. Romero, D. Tzionas, and M. J. Black (2022); O. Taheri, N. Ghorbani, M. J. Black, and D. Tzionas (2020); L. Yang, K. Li, X. Zhan, F. Wu, A. Xu, L. Liu, and C. Lu (2022)), which prevents directly combining different data resources to scale up. The limited scale of object annotations causes the method to overfit to biased object shapes and thus bad generalizability of generating grasps on unseen objects.
Humans have the remarkable ability to transfer grasping patterns to different objects and make plausible grasping choices for unseen objects based solely on their geometry. While prior knowledge, such as object category or functionality, can aid in manipulation, it is often not essential for determining a grasp.
We refer to this capability as universal grasp generation. We argue that the key to universally generating hand-object grasps lies in developing an inclusive object shape embedding. This is unattainable if we train object embeddings using datasets with limited quantity and diversity. Although comprehensive 3D annotations are scarce, we have access to large-scale 3D object datasets. In this work, we investigate whether these datasets can be leveraged to enhance hand-object grasp generation by jointly modeling hand and object representations in a latent space. Our model is designed to accommodate partial supervision during training, accepting available data such as hand poses or object shapes. This flexibility significantly broadens the range of datasets that can be utilized for learning. Our model can generate complete 3D grasping scenes, including both the posed hand and object, or a posed hand conditioned on a given object geometry. We believe that this flexible approach to generating hand-object interactions, without relying on domain-specific auxiliary information, represents a significant step toward human-like universal grasp generation.
@@
Following such an intention, we propose our Joint Hand-Object Diffusion (JHOD) model. It follows the latent diffusion model (Rombach et al., 2022) (LDM) to encode and ensemble different modalities into a latent space and then generate plausible samples by the probabilistic denoising process (Ho et al., 2020). To learn from more diverse objects, we construct the model by leveraging the object shape encoding and decoding networks pre-trained on the large-scale generic 3D object shape dataset (Chang et al., 2015). As we aim to derive a universal grasp generation solution, we remove the category-specific information when constructing the latent code for objects and encode the object shapes in a purely geometric fashion.
@@
As the hand articulation and positions are always coupled to the object shape in hand-object grasping, we design the latent diffusion to denoise the latent code from different modalities as a whole. We decouple each modality representation to allow model training by combining data from different resources with heterogeneous annotations. Therefore, our method has independent encoder and decoder networks for different modalities and can optimize the generation of each modality solely. Such a design relieves the limitation of object diversity in hand-object interaction datasets as we could leverage the data from generic object shape datasets to tune the object generation part. The model is trained to generate the hand part regarding the corresponding object shape embedding, so an inclusive object generation improves the corresponding hand grasp generation on diverse object shapes. To achieve disentangled modality representations to boost training, we propose to use asynchronous denoising schedulers for different modalities during noise diffusion and denoising in training.
@@
In conclusion, we have developed a joint hand-object grasp generation model capable of sampling from either the joint distribution of both hand and object or the object-conditioned distribution for hand generation. This model effectively leverages training resources with varying annotations. Through both qualitative and quantitative evaluations, our method demonstrates strong performance in grasp generation for both unconditional and object-conditioned scenarios. Thanks to the posterior generation learned from extensive data resources, our model significantly outperforms others in generating grasps for out-of-domain object shapes, a crucial aspect of universal grasp generation. Given that no existing works have supported the end-to-end generation of both hand and object to form a grasp, we believe our approach is pioneering in the field of universal grasp generation.
@@
With the flexibility to generate both modalities in hand grasp, our model can facilitate many downstream applications. The grasps are good references for generating photo-realistic images with good geometry alignment and significantly fewer artifacts. We first generate grasps in 3D and use them as conditional signals for image generation (Adobe, 2024) can help produce high-quality hand images with challenging gestures and view angles. We demonstrate using grasp from our model as the condition improves grasp alignment on images and relieves artifacts. Some examples of object-conditioned grasp generation, joint generation, and image generation with grasp rendering are presented in Figure˜1. The main contribution of this work is to propose a generative model capable of generating hand grasp in either unconditional or object-conditioned fashions and the corresponding strategy to enhance the performance with limited available 3D hand-object grasp annotations.
@@ 2 Related Works
Hand Grasp Generation. Modern generative models (Ho et al., 2020) are recently introduced into hand grasp reconstruction (Ye et al., 2023a) and generation. Hand grasp reconstruction asks for hand grasp rendering aligned with images while hand grasp generation asks for hand grasp visually or physically plausible. GrabNet (Taheri et al., 2020) is a commonly used method for hand grasp generation but it can hardly be generalized to diverse object shapes. HOIDiffusion (Zhang et al., 2024b) uses the 3D hand grasp from GrabNet and an image diffusion model to generate hand-object interaction images. AffordanceDiffusion (Ye et al., 2023b) also studies the generation of hand-object interaction for image synthesis. Instead of generating hand-object interaction images or videos, we focus on improving hand-object interaction in 3D space. Many recent works are based on diffusion models (Ye et al., 2024; Lu et al., 2023; Christen et al., 2024) or physics-based simulators (Zhang et al., 2024a; Luo et al., 2024a; Xu et al., 2023; Luo et al., 2024b). Compared to existing works which typically require a given object (Taheri et al., 2020; Zhang et al., 2024b; Ye et al., 2023b; Zhou et al., 2024) shape as input, we desire generating a hand grasp over a given object or generating both the object and hand to form a plausible grasp.
Multi-Modal Generation. With the rise of diffusion models, multi-modal generation has been extended in many areas, such as image+text generation (Bao et al., 2023) and audio+pixel generation (Ruan et al., 2023). Our model can jointly generate the hand and object in a grasp, making another type of multi-modal generation. A concurrent work UGG (Lu et al., 2023) studies to generate both hand and object to form a grasp but it uses the ShadowHand parametric model to leverage the large-scale synthetic ShadowHand grasping datasets (Wang et al., 2023) and the final results are optimization-based instead of directly from the generative model. On the other hand, another concurrent work G-HOP (Ye et al., 2024) builds a multi-modal hand-object grasp prior by encoding object shape and hand poses into a unified latent space but a language prompt is required to provide necessary conditional information. In this work, we focus on a dual-modal latent diffusion model to generate the modalities of a 3D object and hand at the same time without requiring optimization or auxiliary conditions. The hand grasp is purely geometric-based. It allows more flexibility for downstream tasks, such as generating photo-realistic images by using the rendering of generated 3D hand-object interaction as a condition.
Figure 2: The illustration of our proposed Joint Hand-Object Diffusion (JHOD). We present the main modules of the model while removing the secondary modules for simplicity. The optional object condition can turn the generation into object-conditioned.
@@ 3 Method
In  Section˜3.1, we first provide the formal problem formulation. We then review the related preliminary knowledge for methodology and data representation in Section˜3.2. Finally, we introduced our proposed method in Section˜3.3.
@@ 3.1 Problem Formulation
An instance of hand-object grasp consists of a posed hand, denoted as $\mathbf{H}$, and a posed object, denoted as $\mathbf{O}$. To generate a hand-object grasp unconditionally, we model the joint distribution of hand and object:
$\vskip-1.00006pt\{\mathbf{O},\mathbf{H}\}\sim\mathcal{D}_{\Phi}.\vskip-1.00006pt$ (1)
Also, in the community of visual perception, animation, and robotics, people are interested in generating hand grasp conditioned on certain objects, which can be formulated as
$\vskip-1.00006pt\{\mathbf{H}\}\sim\mathcal{D}_{\Phi}|_{\mathbf{O}_{g}},\vskip-1.00006pt$ (2)
where $\mathbf{O}_{g}$ is a provided object shape. In the previous works, the unconditional generation has been under-explored. The methods for conditional hand grasp generation are usually required with no flexibility for object shape generation. They suffer from bad robustness to different object shapes. This is because the existing datasets with 3D annotations of both modalities are notoriously limited in covered object shapes due to the expansiveness of scanning the object shape and capturing the corresponding hand poses. These methods’ limitation is underestimated because the object shapes used in training and testing in existing datasets share similar scales and geometries. In this work, we aim to solve both unconditional and object-conditioned grasp generation by a single model. We also wish to enhance the robustness of object shapes by deriving a more inclusive object embedding which also improves the generalizability of the hand part generation.
@@ 3.2.1 Latent Diffusion
Our method follows the latent diffusion models (LDM) (Rombach et al., 2022; Sohl-Dickstein et al., 2015) to generate samples in two modality spaces, i.e., articulable hands, and rigid objects. LDM diffuses and denoises in a latent space instead of on the raw data representations.
@@
where $\alpha_{t}\in(0,1)$ are constant hyper-parameters determined by the time step $t$ and a noise scheduler. Though the typical diffusion model (Ho et al., 2020) trains a denoiser to predict the noise added per time step during the diffusion stage, there is another line of diffusion models (Ramesh et al., 2022; Tevet et al., 2022) that learns to recover the clean data sample directly. The loss turns to
@@ 3.2.2 Modality Representations
Hand. We follow the MANO (Romero et al., 2022) parametric model to represent hands. Compared to another parametric model ShadowHand (6), MANO is designed for animating non-rigid real human hands instead of robotic hands thus a better fit for animation and photo-realistic generalization tasks. However, its higher complexity causes difficulty in grasping data synthesis. Therefore, there is no MANO-based hand grasping dataset as large-scale as the synthetic ShadowHand-based datasets (Wang et al., 2023) yet. This prevents research on the joint generation of MANO hands and objects in a straightforward supervised learning fashion. In this work, we directly use MANO parameters as joint rotation angles, i.e. $\theta\in\mathbb{R}^{48}$, and the translation of hands $\mathbf{t}_{h}\in\mathbb{R}^{3}$. We use constant mean shape parameters $\beta_{0}=\mathbf{0}^{10}$ during linear blend skinning. The parameters of hands are concatenated and encoded by a hand encoder $\mathcal{E}_{H}$ to be the hand latent code $\mathbf{y}_{H}\in\mathbb{R}^{51}$.
Object. For object representation, we use the point cloud latent code as $\mathbf{y}_{O}=\mathbf{h}\in\mathbb{R}^{N\times 4}$, where $N$ is the number of points. As the grasp is modeled in the object-centric frame, no translation or orientation is required in the object pose and shape representation. This convention is aligned with the definition of data in many hand-object grasp datasets, such as OakInk (Yang et al., 2022).
@@ 3.2.3 Object Generation
The limited amount and high category-specific bias of the objects in hand-object interaction (HOI) datasets prevent learning generating grasps robust and generalizable to diverse object shapes. When training on these datasets only, the learned object embedding is usually biased and overfit. By leveraging the large-scale object-only 3D datasets, we can derive a more robust object encoding for the grasp generation than learning from the HOI datasets solely. For the object encoding part, we borrow the point-cloud-based object shape generation method LION (Vahdat et al., 2022). Existing HOI generation methods can only learn object shapes from HOI datasets, including only dozens of objects, while LION learns from a much larger basis, i.e., more than 50,000 objects in ShapeNet (Chang et al., 2015). LION generates objects in two stages. In the first step, it derives a global latent code $\mathbf{z}^{G}_{0}\in\mathbb{R}^{D_{G}}$ from a posterior distribution $q_{\phi}(\mathbf{z}^{G}_{0}|\mathbf{P})$ where $\mathbf{P}\in\mathbb{R}^{u\times 3}$ is the object point cloud. Then it samples a point cloud latent $\mathbf{h}_{0}$ from a posterior $q_{\mathcal{E}_{O}}(\mathbf{h}_{0}|\mathbf{P},\mathbf{z}^{G}_{0})$. LION generates the object shape by a reverse process from sampled latent code
$\vskip 0.0ptp_{\mathcal{E}_{O},\psi,\gamma}(\mathbf{P},\mathbf{h}_{0},\mathbf{z}^{G}_{0})=p_{\mathcal{X}_{O}}(\mathbf{P}|\mathbf{h}_{0},\mathbf{z}^{G}_{0})p_{\psi}(\mathbf{h}_{0}|\mathbf{z}^{G}_{0})p_{\gamma}(\mathbf{z}^{G}_{0}),\vskip 0.0pt$ (5)
@@
$\vskip 0.0ptp_{\mathcal{E}_{O},\psi^{\prime}}(\mathbf{P},\mathbf{h}_{0})=p_{\mathcal{X}_{O}}(\mathbf{P}|\mathbf{h}_{0})p_{\psi^{\prime}}(\mathbf{h}_{0}),\vskip 0.0pt$ (6)
where $\psi^{\prime}$ is fine-tuned from $\psi$ by replacing the global prior $\gamma$ with a standard Gaussian. Thanks to the pre-training from large-scale object datasets (Chang et al., 2015), we could use the weights of $\mathcal{E}_{O}$ and $\mathcal{X}_{O}$ from LION directly. By freezing the encoder and decoder, we fine-tune the object prior distribution with the object shapes from HOI datasets.
3.3 Joint Hand-Object Diffusion (JHOD)
We now introduce our proposed Joint Hand-Object Diffusion (JHOD). There are two main goals. First, we would like to use the training data with heterogeneous annotations so that we can go beyond the limited object shapes in HOI datasets. Second, we want a single model capable of object-conditioned hand grasp generation and unconditional hand-object joint generation. We present the overall illustration and key designs of our proposed method in Figure˜2.